OpenStudy (superflytyguy):

If a function, f(x) is shifted to the left four units, what will the transformed function look like?

OpenStudy (utterly_confuzzled):

For an equation like f(x)=x^2 it would become f(x)=(x+4)^2 or y=f(x+4)

OpenStudy (anonymous):

F(x + 4) + c

OpenStudy (superflytyguy):

Thank you!

OpenStudy (superflytyguy):

Both of you!

OpenStudy (anonymous):

you're welcome!

OpenStudy (utterly_confuzzled):

You're welcome!!